64 gram methane are mixed with 64 gram of oxygen and burnt,then the mass of carbondioxide obtained is?
1
Expert's answer
2019-05-28T07:28:44-0400

The Stoichiometric equation for the reaction is

CH4 + 2O2 = CO2 + 2H2O

Moles of CH4  = 64/16 = 4

Moles of O2 = 64/32 = 2

For every 2 moles of oxygen 1 mole of methane is required thus oxygen is the limiting reactant here.

For 2 moles of oxygen 1 mole of CO2 is obtained

Thus 2 moles of Oxygen will give 1 mole of carbon dioxide thus giving 44g.